OpenStudy (anonymous):

Teacher wont help me. Someone please explain how to do this! Solve for X: 2-4/x^2=-2/x

OpenStudy (anonymous):

Those are fractions!

OpenStudy (anonymous):

\[2-\frac{4}{x^2}=-\frac{2}{x}\] like that?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

since \(x\) cannot be zero, a good first step is to multiply both sides by \(x^2\) and rewrite as \[2x^2-4=-2x\]

OpenStudy (anonymous):

then add \(2x\) to both sides and get \[2x^2+2x-4=0\] divide by 2 to get \[x^2+x-2=0\] and then either factor or use the quadratic formula

OpenStudy (anonymous):

if you don't understand any step let me know